How many high energy molecules are made or used in converting citrate to glucose. Please show the steps if you can!
Answer #1

Answer: 15 molecules of high emergy compiunds are produced by production of 1 glucose molecule from 1 citrate.

Making of a glucose molecule is termed gluconeogenesis. The path of making glucose from citrate is as follows:

NAD+ NADH + H+ NAD+ NADH + H+ Citrate Isocitrate a-ketoglutarate Succinyl COA 4CO2 CO2 GDP + Pi GTP Succinate TCA cycle react

GTP GDP Oxaloacetate Phosphoen olpyruvate 9 2-phosphoglycerate 10 3-phosphoglycerate 8 ADP 11 ATP Oxaloacetate to glucose rea

Note that since glucose is a 6 carbon molecule, and oxaloacetate is 4 carbon molecule which eventually decarboxylates to phosphoenolpyruvate, a 3 carbon molecule, all the molecules up to reaction 12 is indeed needed in 2 moles for production of 1 mole of glucose. That is why, prior reactants are prefixed with a half (1/2), indicating that the molecule will only produce half a mole glucose. Our approach is to first calculate how many high-energy compounds (ATP, GTP) produced or released during the entire process.

From reactions, 1-16,

ATP/GTP produced= 2

ATP/GTP released= 1

Net ATP/GTP= +1 ( 1 ATP/GTP produced from the entire process)

But NADH and FADH2 are reducing equivalents which can produce ATP through Electron Transport System. NADH generates about 2.5 ATP per molecule. Whereas FADH2 producesa about 1.5 ATP per molecule. So, let's count NADH and FADH2 MOLECULES;

NADH produced: 3

NADH expense=1

Net NADH= 3-1= +2 or 2 NADH molecules are produced

FADH2 produced=1

FADH2 expense: none

Net FADH2= +1 or Net 1 FADH2 molecule is generated in the entire process.

Converting Net NADH to ATP= 2*2.5= 5 ATP

Converting Net FADH2 to ATP= 1*1.5= 1.5 ATP

Hence, Total ATP produced = 5 (from NADH) + 1.5 (from FADH2) + 1 (from ATP/GTP)= 7.5 ATP

Thus, 7.5 molecules of ATP is generated to produce half glucose molecule.

Hence, 1 molecule of glucose generation through this route will produce 7.5*2= 15 molecules of ATP